OPTICAL STARK SPECTROSCOPY OF THE B^{1}A′′-X^{1}A′ of CuOH

\underline{FANG WANG} AND TIMOTHY C. STEIMLE, Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287.

Copper hydroxide, CuOH, is one of the most thoroughly studied ^{a} transition metal containing polyatomic molecules with the most recent investigation being the dispersed fluorescence and radiative lifetime study. ^{b} The pure rotational spectrum was recorded and analyzed by Saito′s group some time ago.^{c} Recently a high level ab initio calculation predicted ground state permanent electric dipole moments, μ_{el}, ranging in value from 5.52D to 3.98D depending upon methodology used to treat electron correlation. Here we report on the first molecular beam study of the B^{1}A′′-X^{1}A′ band. The spectrum was recorded field-free and in the presence of a static electric field of up to 4000 V/cm. The hyperfine splitting is observed and analyzed for the first time. The determined μ_{el} values are used to evaluate the computational methodologies and are compared with those of other copper containing molecules.

aC.N.Jarman, W.T.M.L Fernando, and P.F. Bernath J.Mol.Spectrosc.. 144 286, 1990.

bC.Tao, C. Mukarakate, and S. A. Reid Chem.Phys.Lett.. 449 282, 2007.

cC.J. Whitham, H. Ozeki, and S. Saito J.Chem.Phys.. 112 641, 2000.